1. Prior service cost 2. Prior service cost amortization Service cost 3. Interest cost 5. Pension expense 4.Sachs Brandss defined benefit pension plan specifies annual retirement benefits equal to 1.6% × service years x final years
Answer #1

Answer Page: no ① 0 Price device est PBO with ouf amendment cosf - 1-6%. X service years x final years sclery I PUAP (67., 18Page: No @ Prior service cost = 221553 - 202563 = $18990 Pui Prior service cost = $18990 © Prior service cost amortization :Page No ③ = 1.454. X1 X $250,000 X 10.82760 0.33051 - 15657 Lii Service cosf = $15657 © Einterest cook Interesti coof - PBO w